OpenStudy (elise_a18):

Sam is observing the velocity of a car at different times. After two hours, the velocity of the car is 54 km/h. After four hours, the velocity of the car is 58 km/h. Part A: Write an equation in two variables in the standard form that can be used to describe the velocity of the car at different times. Show your work and define the variables used.

undeadknight26 (undeadknight26):

The answer is always Nutella.

undeadknight26 (undeadknight26):

Aka its all over google Elise hun ._. http://openstudy.com/study#/updates/53776427e4b0e5430794e16c http://openstudy.com/study#/updates/52ae1d43e4b0f72fdcf7b2ed

OpenStudy (elise_a18):

I have a function I just need help putting it into standard form f(x)= 52(1.93)^2

OpenStudy (elise_a18):

@AllTehMaffs @Mashy @mathhelper @ilovemath

OpenStudy (anonymous):

How'd you get that? And either of those links that are posted are very helpful as a step by step through the problem.

OpenStudy (elise_a18):

every2 2 hrs it increases at a rate of 93% and 52 is the y-int

OpenStudy (elise_a18):

That way it is a quadratic function and can be put in standard form

OpenStudy (anonymous):

Neither of those statements are actually true unfortunately. How did you get the 93% ? If this is our graph If the velocity increases by 4km/h every 2 hours, then 2 hours before it was 54km/h it was 54km/h-4km/h at t=0, and t=0 is where our axes cross

OpenStudy (anonymous):

You don't need to use a quadratic equation - it actually specifically won't fit this graph because the velocity increases as a straight line. If you look at the other threads that were posted, they go over that the two variables you need are the acceleration and the time, and you can find the acceleration by \[a = \frac{\Delta v}{\Delta t}= \frac{v_2-v_1}{t_2-t_1}\] the change in velocity over the change in time - it's exactly the same form as the slope of a line in the form \[y = mx + b\] with m being analogous to a, t being analogous to x, and y being analogous to v, so \[v = at + b\] your two ordered pairs are the information you need to find the acceleration (aka, the slope of the line)

OpenStudy (anonymous):

I have to run, sorry. I'll see if you still need help when you get back.

OpenStudy (elise_a18):

@Loser66

OpenStudy (elise_a18):

@mathmath333 @Mashy @HelpBlahBlahBlah @KyanTheDoodle @omarbirjas

OpenStudy (elise_a18):

@dan815 @dtan5457

OpenStudy (anonymous):

Are you still stuck on a part of it?